OpenStudy (brittany16):

Solve for the variable. m - 4 = 18

OpenStudy (anonymous):

m - 4 = 18 m=14

jimthompson5910 (jim_thompson5910):

m - 4 = 18 m - 4+4 = 18 + 4 m = 22